You just won the lottery. You can take your Rs.10, 00,000 in a lump- sum today, or you can receive Rs. 100,000 at the end of every year for 12 years. You can invest your money 3%per annum. Ignoring all tax consideration, which would you prefer? But if you receive Rs.50,000at the end of every six-month as semi- annual compounding system, what could be the difference in your decision?







1
Expert's answer
2022-02-28T11:22:25-0500

The future value of a series of annuity payments (FV) is calculated by the formula


"FV=A(\frac{(1+r)^n)-1}{r}"

A=100 000

r=0.03

n=12

"FV=100 000(\frac{(1+0.03)^{12})-1}{0.03}=1419 202.96"


A=100 000

r=0.015

n=24

"FV=100 000(\frac{(1+0.015)^{24})-1}{0.015}=1 304 121.14"


A total of 100,000 personal goods and invested in 3% of goods in a 12-year lease